The present value of a loan of R12 000 due in five years’ time, at a simple discount rate of 15% per annum is


[1] R24 136,29.

[2] R48 000,00.

[3] R3 000,00.

[4] R5 324,46.

[5] none of the above.


1
Expert's answer
2021-07-06T04:35:04-0400

interest =Principal x interest x period 

Principal =RM 12000

interest rate =15%

Period = 5 years 

Interest =12000 × 0.15 x 5=9000

Present value =Principal amount"-" interest 

Present value=12000-9000=3000

Present value =R 3000

The present value of a loan of R12 000 due in five years’ time, at a simple discount rate of 15% per annum is R3000 

[3] R3 000,00.